OpenStudy (yanasidlinskiy):

Check this!!:) Thanx!! Find the probability. An IRS auditor randomly selects 3 tax returns (without replacement) from 47 returns of which 5 contain errors. What is the probability that she selects none of those containing errors? A. 0.0012 B. 0.0006 C. 0.708 D. 0.7136 This is D. Correct?

OpenStudy (yanasidlinskiy):

A lot of people are confusing me. Am I right?

OpenStudy (nurali):

Total tax returns = 47 of which Tax returns containing errors = 5 Tax returns not containing errors = 47 - 5 = 42 Required probability = 42C3*5C0/47C3 = 11480*1/16215 = 0.707986 or 70.7986%

OpenStudy (yanasidlinskiy):

So you're for sure it's C??? I really don'e want to get this wrong:P

OpenStudy (nurali):

Ya

OpenStudy (yanasidlinskiy):

Btw, thanx sooo much!!!:)

OpenStudy (nurali):

Anytime.

OpenStudy (yanasidlinskiy):

Can you answer another question?

OpenStudy (yanasidlinskiy):

In a few minutes though....Hang on..